# softMathematical software

### v1

@@ 2.2.3. Structural semi-graphoids
The supermodular difference expression for $ij|K\in{\mathcal{K}}(N)$ is a special case of it:
@@
The frame $\mathfrak{st}$ is closed under copying, marginalization, intersection, and duality.
Proof 2.12.
Given a bijection $\varphi:N\to M$ and a set function $m:{\mathcal{P}}(N)\to{\mathbb{R}}$, one has $\Delta m(ij|K)=\Delta(m\circ\varphi_{-1})\,\varphi(ij|K)$ for any $ij|K\in{\mathcal{K}}(N)$, where $\varphi$ is naturally extended to the power set of $N$ and ${\mathcal{K}}(N)$. Hence, provided that $m$ is a supermodular function on ${\mathcal{P}}(N)$, the composition $m\circ\varphi_{-1}$ is a supermodular function on ${\mathcal{P}}(M)$ and the $\varphi$-copy of ${\mathcal{M}}^{m}$ is ${\mathcal{M}}^{m\circ\varphi_{-1}}$.
@@
Basic facts are that $\mathfrak{pr}\subseteq\mathfrak{st}\subseteq\mathfrak{sg}$ and $\mathfrak{sg}(N)=\mathfrak{st}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$; for $\mathfrak{pr}\subseteq\mathfrak{st}$ see [29, § 2.3.4]. The strict inclusion $\mathfrak{st}(N)\subset\mathfrak{sg}(N)$ in case $|N|=4$ is evidenced by the model :=${\mathcal{M}}:=\{\,ab|c,\,ac|d,\,ad|b\,\}$ from Example 2.8, because the property (E:1) mentioned there holds for structural semi-graphoids.
The strict inclusion $\mathfrak{pr}(N)\subset\mathfrak{st}(N)$ in case $|N|=4$ is evidenced by the dual model :=${\mathcal{M}}^{\star}:=\{\,ab|\emptyset,\,ab|c,\,ab|d,\,cd|ab\,\}$ from Example 2.5, which is induced by the following supermodular function $m$ on $N=\{a,b,c,d\}$:
@@
Remarks.  This is to remind alternative views.
@@
    The concept of a structural semi-graphoid can be viewed as a reduced version of the concept of a semi-matroid from [24, § 2]. The difference is that semi-matroids involve also functional dependence statements, interpreted as special CI statements $i\perp\!\!\!\perp i\,|\,K$ for $i\in N$, $K\subseteq N\setminus\{i\}$. But otherwise semi-matroids are defined analogously: they are induced by polymatroids through vanishing of the respective submodular difference expressions.
@@
2.3. Lattices of models and closure operations
Given a CI frame $\mathfrak{f}$ closed under intersection and a fixed variable set $N$, the poset $(\mathfrak{f}(N),\subseteq)$ appears to be a finite lattice. Such a lattice can be characterized in two methodologically different ways:
  • in terms of irreducible models (= an inner description in the sense that the characterization only depends on the lattice itself),
In this section we recall relevant concepts from lattice theory [1].
An elementary observation is that any Moore family $({\mathcal{F}},\subseteq)$, equipped with the set inclusion relation, is a complete lattice [1, Theorem 2 in § V.1]. [^2] This is a universal construction in the sense that any lattice is isomorphic to a lattice based on a Moore family and any finite lattice is isomorphic to a finite Moore family lattice. These conclusions can be derived from observations in [1, § V.9]. In our treatment the intended setting is $X:={\mathcal{K}}(N)$ for a finite variable set $N$. Thus, we have the following observation.
Corollary 1.
If a CI frame $\mathfrak{f}$ is closed under intersection then, for each set $N$, the family $(\mathfrak{f}(N),\subseteq)$ is a finite lattice in which infimum means intersection. If $\mathfrak{f}$ is also closed under copying and $|N|=|M|$ then $(\mathfrak{f}(N),\subseteq)$ and $(\mathfrak{f}(M),\subseteq)$ are order-isomorphic and permutations of $N$ yield automorphisms of $(\mathfrak{f}(N),\subseteq)$.
Proof 2.13.
As $\mathfrak{f}(N)$ is a Moore family, this is evident from what was said before.
Let $({\mathcal{L}},\preceq)$ be a (finite) lattice with the greatest element denoted by 1 and the meet/infimum operator by $\inf$. Let $\prec$  denote the strict order relation. Given $\lambda,\tau\in{\mathcal{L}}$, one says that $\tau$ covers $\lambda$ and writes $\lambda\prec\!\!\!\cdot\,\,\,\tau$ if $\lambda\prec\tau$ and there is no $\kappa\in{\mathcal{L}}$ with $\lambda\prec\kappa\prec\tau\/$. A coatom of ${\mathcal{L}}$ is a sub-maximal element in ${\mathcal{L}}$, that is, $\tau\in{\mathcal{L}}$ satisfying $\tau\prec\!\!\!\cdot\,\,\,\mbox{\bf 1}$. An element $\lambda\in{\mathcal{L}}$ is called (meet) irreducible if $\lambda\neq\inf\,\{\,\tau\in{\mathcal{L}}\,:\ \lambda\prec\tau\,\}$. [^3] Note that 1 is not irreducible by this definition: it is the infimum of the empty set. A well-known fact is that the set ${\mathcal{I}}$ of (meet) irreducible elements in a finite lattice $({\mathcal{L}},\preceq)$ is the least set determining all elements of ${\mathcal{L}}$ by means of the infimum operation in the following sense:
The coatoms are standard examples of irreducible elements in ${\mathcal{L}}$. In general, a lattice can have other irreducible elements. A lattice $({\mathcal{L}},\preceq)$ is called coatomistic if every element of ${\mathcal{L}}$ is the infimum of a set of coatoms, which is another way of saying that there is no irreducible element in ${\mathcal{L}}$ except the coatoms.
Example 2.14.
Put $X:=\{x,y,z,w\}$ and consider a Moore family
:=${\mathcal{F}}:=\{\,\,\{x,y,z,w\},\{x,y,z\},\{x,y,w\},\{x,y\},\{x\}\,\,\}\,,$
Figure 1. A Moore family $\{\,\{x,y,z,w\},\{x,y,z\},\{x,y,w\},\{x,y\},\{x\}\,\}$.
Thus, the first way to characterize a finite lattice $({\mathcal{F}},\subseteq)$, where ${\mathcal{F}}$ is a Moore family of subsets of $X$, is to give the complete list of irreducible elements in ${\mathcal{F}}$. We now apply this to the case $X:={\mathcal{K}}(N)$ for a finite variable set $N$ and ${\mathcal{F}}:=\mathfrak{f}(N)$ being a suitable family (of CI models) over $N$.
Corollary 2.
If a CI frame $\mathfrak{f}$ is closed under copying and intersection then, for each variable set $N$, the irreducible sets in $(\mathfrak{f}(N),\subseteq)$ (= irreducible models) fall into permutational equivalence classes (= relative to permutations of $N$).
Proof 2.15.
By Corollary 1, any permutation $\pi:N\to N$ defines an order automorphism of $(\mathfrak{f}(N),\subseteq)$, namely the natural extension of $\pi$ to ${\mathcal{K}}(N)$. Such a transformation preserves the irreducibility of an element of $\mathfrak{f}(N)$.
Example 2.16.
(lattice of semi-graphoids over 3 variables)
Let us take $N:=\{a,b,c\}$, $X:={\mathcal{K}}(N)$, and put ${\mathcal{F}}:=\mathfrak{sg}\,(N)$. The lattice $(\mathfrak{sg}\,(N),\subseteq)$ has 22 elements in this case and they break into 10 permutational equivalence classes. There are 5 (meet) irreducible elements in $\mathfrak{sg}\,(\{a,b,c\})$ which fall down into 3 permutational types:
  • $\{\,ab|\emptyset,\,ac|\emptyset,\,bc|\emptyset\,\}$,
  • $\{\,ab|c,\,ac|b,\,bc|a\,\}$,
A kind of dual description of a finite lattice is based on the following concept. A closure operation on subsets of $X$ is a mapping $\mbox{\rm cl}:{\mathcal{P}}(X)\to{\mathcal{P}}(X)$ which is
  • extensive: $Y\subseteq\mbox{\rm cl}\,(Y)$ for $Y\subseteq X$,
A set $Y\subseteq X$ is then called closed with respect to cl if $Y=\mbox{\rm cl}\,(Y)$. A basic fact is that the collection of subsets of $X$ closed with respect to cl is a Moore family and any Moore family can be defined in this way [1, Theorem 1 in § V.1]. Specifically, given a Moore family ${\mathcal{F}}$ of subsets of $X$, the formula
defines a closure operation on subsets of $X$ having ${\mathcal{F}}$ as the collection of closed sets with respect to $\mbox{\rm cl}_{{\mathcal{F}}}$; see [13, Theorem 1 in § 0.3]. Thus, for any $Y\subseteq X$, the closure $\mbox{\rm cl}_{{\mathcal{F}}}(Y)$ of $Y$ relative to ${\mathcal{F}}$ is the smallest element of ${\mathcal{F}}$ containing $Y$.
Example 2.17.
The closure operation can also be interpreted as a functional dependence relation on subsets of $X$, as explained in [19]. This is a binary relation $\to$ on ${\mathcal{P}}(X)$ satisfying three axioms (for $Y,Z,W\subseteq X$):
  • reflexivity: $Z\subseteq Y~{}\Rightarrow~{}Y\to Z$,
  • transitivity: $[\,Y\to Z~{}\&~{}Z\to W\,]~{}\Rightarrow~{}Y\to W$,
  • union: $[\,Y\to Z~{}\&~{}Y\to W\,]~{}\Rightarrow~{}Y\to Z\cup W$.
Together, these axioms are equivalent to the well-known Armstrong axioms for functional dependence relations in database theory. Given a closure operation cl on subsets of $X$, define $Y\to Z$ as $Z\subseteq\mbox{\rm cl}\,(Y)$. Given a functional dependence relation $\to$, put $\mbox{\rm cl}\,(Y):=\bigcup\,\{\,Z\subseteq X\,:\ Y\to Z\,\}$, which closure operation induces the original functional dependence relation $\to$.
Example 2.18.
[…]
The second way to characterize a finite lattice $({\mathcal{F}},\subseteq)$, where ${\mathcal{F}}$ is a Moore family of subsets of $X$, is to give a suitable generator of the respective functional dependence relation $\to$. Before explaining the details we wish the reader to realize/observe that there is a straightforward relationship (= correspondence) between the functional dependence relations $\to$ and Moore families ${\mathcal{F}}$ [^4] A one-to-one assignment between a Moore family and a functional dependence relation has already been established through their shared closure operation. We now re-interpret this as a kind of duality relationship and avoid the concept of a closure operation.
Specifically, any element $(Y,Z)$ of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ can be interpreted as a formal implication $Y\to Z$, meaning that the set $Y$ of antecedents implies (all elements of) the set $Z$ of consequents. One can introduce a binary relation between elements of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and ${\mathcal{P}}(X)$ as follows:
$(Y\to Z)\odot F\quad:=\quad[\,\,Y\setminus F\neq\emptyset~{}~{}\mbox{or}~{}~{}Z\subseteq F\,\,],$
that is, $[\,Y\subseteq F~{}\Rightarrow~{}Z\subseteq F\,]$, meaning that $Y\to Z$ is valid for $F$.
The next step is to define the so-called Galois connections between subsets of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and subsets of ${\mathcal{P}}(X)$:
Well-known classic results in lattice theory on Galois connections [1, § V.7], see also specifically [19, Theorem], say the following:
  • the mapping ${\mathbb{G}}\mapsto{\mathbb{G}}^{\triangleright\triangleleft}$ is a closure operation on subsets of (a much larger set) ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and ${\mathbb{G}}^{\triangleright\triangleleft}$ is then the smallest functional dependence relation (on subsets of $X$) containing ${\mathbb{G}}$,
  • the mapping ${\mathcal{F}}\mapsto{{\mathcal{F}}}^{\triangleleft\triangleright}$ is a closure operation on subsets of ${\mathcal{P}}(X)$ and ${{\mathcal{F}}}^{\triangleleft\triangleright}$ is then the smallest Moore family (of subsets of $X$) containing ${\mathcal{F}}$,
Example 2.19.
Consider $X:=\{x,y,z,w\}$ and put
${\mathbb{G}}:=\{~{}\emptyset\to\{x\},~{}\{z\}\to\{y\},~{}\{x,w\}\to\{x,y,w\}~{}\}\,.$
Then ${\mathbb{G}}^{\triangleright}$ equals to the family ${\mathcal{F}}$ from Example 2.14. Hence, ${\mathbb{G}}$ is one of several possible implicational generators of ${\mathcal{F}}$, a minimal one in sense of inclusion.
To introduce a prominent implicational generator one needs specific concepts from (theory of) formal concept analysis; see [12, Chapter  3] or [34, § 1.3.2]. [^5] In general, there is no unique smallest implicational generator is sense of inclusion.
Definition 2.20 (pseudo-closed sets, canonical basis).
We say that $Q\subseteq X$ is pseudo-closed relative to a Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$ if
  • $Q\not\in{\mathcal{F}}$ (= $Q$ is not closed),
The recursive definition of pseudo-closedness of a set $Q$ is well-formed as it depends only on the previously defined pseudo-closedness of strict subsets of $Q$. The canonical implicational basis, alternatively named Duquenne–Guigues basis, was introduced already in the 1980s [15]. We have also implemented an algorithm to compute the canonical basis ${\mathbb{G}}_{\mathcal{F}}$ for a given Moore family ${\mathcal{F}}$; its idea is described in Appendix A.
Example 2.21.
$\{~{}\emptyset\to\{x\},\,\{z\}\to\{y\},\,\{w\}\to\{y\}~{}\}\,.$
The well-known result [12, Theorem 7 in § 3.2] says $({\mathbb{G}}_{{\mathcal{F}}})^{\triangleright}={\mathcal{F}}$ for any Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$. In fact, ${\mathbb{G}}_{{\mathcal{F}}}$ is a non-redundant implicational generator of ${\mathcal{F}}$ in sense that the removal of any of its elements leads to a non-generator.
Let us introduce a specific terminology concerning formal implications. Given a Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$, a valid implication for ${\mathcal{F}}$ is $(Y\to Z)\in{\mathcal{F}}^{\triangleleft}$. An implication $Y\to Z$ is perfect (for ${\mathcal{F}}$) if $Y\not\in{\mathcal{F}}$ but every strict subset of $Y$ is closed, that is, $Y\in({\mathcal{P}}(X)\setminus{\mathcal{F}})_{\min}$, and $Z=\mbox{\rm cl}_{{\mathcal{F}}}(Y)\setminus Y$. The implication is canonical if it belongs to the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$.
Definition 2.22 (implicatively perfect Moore family).
It makes no problem to deduce from non-redundancy of the canonical basis that an equivalent condition is that every pseudo-closed set is an inclusion-minimal non-closed set, that is, $\wp({\mathcal{F}})=({\mathcal{P}}(X)\setminus{\mathcal{F}})_{\min}$. [^6] The concept from Definition 2.22 can be viewed as a counter-part of the concept of a coatomistic lattice in the context of implicational description of the lattice.
Example 2.23.
(semi-graphoids over 3 variables revisited)
Consider again the situation from Example 2.16: $N:=\{a,b,c\}$, $X:={\mathcal{K}}(N)$, and ${\mathcal{F}}:=\mathfrak{sg}\,(N)$. There are six pseudo-closed sets relative to ${\mathcal{F}}$ belonging to one permutational type: the set $\{\,ab|\emptyset,\,ac|b\,\}$ and five of its permutational copies. Each of these sets is minimal in ${\mathcal{P}}(X)\setminus{\mathcal{F}}$; thus, ${\mathcal{F}}$ is implicatively perfect. The canonical basis ${\mathbb{G}}_{{\mathcal{F}}}$ consists of permutational copies of the implication
$\{\,ab|\emptyset,\,ac|b\,\}\to\{\,ac|\emptyset,\,ab|c\,\}\,.$
Thus, the characterization of ${\mathcal{F}}$ in terms of its canonical implicational basis is equivalent to the definition of a semi-graphoid from Section 2.2.2.
The fact that the elements of the canonical basis for semi-graphoids fall down into permutational types is not a surprise because we have the following.
Corollary 3.
If a CI frame $\mathfrak{f}$ is closed under copying and intersection then, for each set $N$, the valid implications $Y\to Z$ and the elements of the canonical implicational basis for ${\mathcal{F}}:=\mathfrak{f}(N)$ fall into permutational equivalence classes.
Proof 2.24.
The arguments are analogous to those in Corollary 2. Any permutation $\pi:N\to N$ defines a simultaneous order automorphism of the posets $(\mathfrak{f}(N),\subseteq)$ and $({\mathcal{P}}({\mathcal{K}}(N)),\subseteq)$. As such an order automorphism of $({\mathcal{F}},\subseteq)$ commutes with the closure operation $\mbox{\rm cl}_{{\mathcal{F}}}$, it preserves the validity of implications.
In the sequel, we freely switch between the description of a CI frame $\mathfrak{f}$ in terms of the families $\mathfrak{f}(N)$ and in terms of the corresponding closure operations.
Definition 2.25 (frame-based closure of a CI model).
Consider a CI frame $\mathfrak{f}$ closed under marginalization, intersection, and satisfying $\mathfrak{f}(N)\subseteq\mathfrak{f}(M)$ if $N\subseteq M$ [^7] As ${\mathcal{K}}(N)\subseteq{\mathcal{K}}(M)$ if $N\subseteq M$, the inclusion $\mathfrak{f}(N)\subseteq\mathfrak{f}(M)$ of set systems makes sense. Given a CI model ${\mathcal{R}}\subseteq{\mathcal{K}}(N)$ over a variable set $N$, the closure of ${\mathcal{R}}$ determined by $\mathfrak{f}(N)$ will shortly be denoted by $\mathfrak{f}({\mathcal{R}}):=\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{R}})$.
The assumptions in Definition 2.25 imply that the definition of $\mathfrak{f}$-closure does not actually depend on $N$: one has $\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{R}})=\mbox{\rm cl}_{\mathfrak{f}(M)}({\mathcal{R}})$ for any $M\supseteq N$. All frames considered in this article satisfy the assumptions of Definition 2.25.
@@ 3. Adhesion and self-adhesion
It has been observed recently in [2] that these concepts can also be defined within a simpler discrete environment of formal CI structures. We follow this link here. An auxiliary concept of consonant models is needed.
@@ 3.1. Self-adhesion
Our basic notion is self-adhesion of a model ${\mathcal{M}}$ is nothing but the adhesion of ${\mathcal{M}}$ with its own $\varphi$-copy, where $\varphi$ is an appropriate bijection. The derived concept of self-adhesivity is relative to a considered CI frame.
@@
  • Let $\mathfrak{f}$ be a CI frame closed under marginalization and copying such that $\mathfrak{f}\subseteq\mathfrak{sg}$. Given ${\mathcal{M}}\in\mathfrak{f}(N)$ and $L\subseteq N$ we say that ${\mathcal{M}}$ is self-adhesive at $L$ relative to $\mathfrak{f}$ if, for every bijection $\varphi:N\to M$ such that $L=N\cap M$ and $\varphi_{\,|L}$ is the identity, an adhesion of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ exists which is in the family $\mathfrak{f}(NM)$:
@@
    Then ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive relative to $\mathfrak{f}$ if it is self-adhesive at any $L\subseteq N$.
  • The symbol $\mathfrak{f}^{\diamond}(N)$ will denote the family of self-adhesive models ${\mathcal{M}}\in\mathfrak{f}(N)$ relative to $\mathfrak{f}$. If $\mathfrak{f}(N)=\mathfrak{f}^{\diamond}(N)$ then we say that $\mathfrak{f}(N)$ is self-adhesive. The frame $\mathfrak{f}$ is self-adhesive if this holds for any variable set $N$.
@@
Note that the assumption $\mathfrak{f}\subseteq\mathfrak{sg}$ is to make the adhesion concept sensible. The other two assumptions (on marginalization and copying) are also needed to define the concept of self-adhesion. Clearly, a model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L\subseteq N$ relative to $\mathfrak{f}$ if and only if $\mathfrak{f}^{\diamond}({\mathcal{M}}|L)={\mathcal{M}}$.
@@
    This is to explain why the self-adhesivity concept is relative to a CI frame. Having two frames $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}$, imagine one finds a model ${\mathcal{M}}\in\mathfrak{f}_{1}(N)$ and a bijection $\varphi:N\to M$ with $L=N\cap M$ such that a self-adhesion ${\mathcal{E}}\in\mathfrak{f}_{2}(NM)$ of ${\mathcal{M}}$ for $\varphi$ exists while there is no such a self-adhesion within $\mathfrak{f}_{1}(NM)$. Then ${\mathcal{M}}$ is self-adhesive at $L$ relative to $\mathfrak{f}_{2}$ but not relative to $\mathfrak{f}_{1}$. This really can happen as Example 5.4 shows (see Section 5.6).
@@
    It follows from the definition that, given ${\mathcal{M}}\in\mathfrak{f}(N)$, the question whether ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ is decidable on basis of the characterization/description of families $\mathfrak{f}(M)$, where $|N|\leq|M|\leq 2\cdot|N|$. Nevertheless, further reduction is possible. For example, it is immediate that any ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L:=N$. Additional reductions, based on further assumptions on the frame, are discussed in Section 3.2.
If the considered frame $\mathfrak{f}$ is closed under intersection then the self-adhesive closure of any model at any set is easy to compute by applying the $\mathfrak{f}$-closure. In particular, the question of self-adhesivity testing can be simplified.
@@
Let $\mathfrak{f}$ be a CI frame closed under marginalization, copying, and intersection such that $\mathfrak{f}\subseteq\mathfrak{sg}$. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $L\subseteq N$, the self-adhesive closure ${\mathcal{R}}=\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$ exists and can be computed as follows. Having fixed a bijection $\varphi:N\to M$ with $L=N\cap M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$, apply the next steps:
@@
    apply the $\mathfrak{f}$-closure to it: ~$\widetilde{{\mathcal{E}}}:=\mathfrak{f}({\mathcal{E}})$,

### v3

@@ 3.2. Structural semi-graphoids
The supermodular difference expression for $ij|K\in{\mathcal{K}}(N)$ is a special case:
@@
$\mathfrak{st}$ is closed under copying, marginalization, intersection, duality, and ascetic extension.
Proof.
Given a bijection $\varphi:N\to M$ and $m:{\mathcal{P}}(N)\to{\mathbb{R}}$, one has $\Delta m(ij|K)=\Delta(m\circ\varphi_{-1})\,\varphi(ij|K)$ for any $ij|K\in{\mathcal{K}}(N)$, where $\varphi$ is naturally extended to the power set of $N$ and ${\mathcal{K}}(N)$. Hence, provided that $m$ is a supermodular function on ${\mathcal{P}}(N)$, the composition $m\circ\varphi_{-1}$ is a supermodular function on ${\mathcal{P}}(M)$ and the $\varphi$-copy of ${\mathcal{M}}^{m}$ is ${\mathcal{M}}^{m\circ\varphi_{-1}}$.
@@
To show that $\mathfrak{st}$ is closed under extension, realize that, if $N\subseteq M$, then any supermodular function $r:{\mathcal{P}}(N)\to{\mathbb{R}}$ can be extended to a supermodular function $m$ on ${\mathcal{P}}(M)$ by the formula $m(T):=r(T\cap N)$ for any $T\subseteq M$. Since $\mathfrak{st}$ is closed under intersection, to show it is closed under ascetic extension it remains to observe that ${\mathcal{K}}(N)\in\mathfrak{st}(M)$. This follows from $\mathfrak{pr}\subseteq\mathfrak{st}$; a specific construction is left to the reader. ∎
Basic facts are that $\mathfrak{pr}\subseteq\mathfrak{st}\subseteq\mathfrak{sg}$ and $\mathfrak{sg}(N)=\mathfrak{st}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$; for $\mathfrak{pr}\subseteq\mathfrak{st}$ see [37, § 2.3.4]. The strict inclusion $\mathfrak{st}(N)\subset\mathfrak{sg}(N)$ in case $|N|=4$ is evidenced by the model :=${\mathcal{M}}:=\{\,ab|c,\,ac|d,\,ad|b\,\}$ from Example 3.2, because the property (E:1) mentioned there holds for structural semi-graphoids.
The strict inclusion $\mathfrak{pr}(N)\subset\mathfrak{st}(N)$ for $|N|=4$ is evidenced by :=${\mathcal{M}}^{\star}:=\{\,ab|\emptyset,\,ab|c,\,ab|d,\,cd|ab\,\}$, which is the dual model in Example 2.6, induced by the next supermodular function $m$ on $N=\{a,b,c,d\}$:
@@
Remarks.  This is to remind alternative views on structural semi-graphoids.
@@
    The concept of a structural semi-graphoid can be viewed as a reduced version of the concept of a semi-matroid from [30, § 2]. The difference is that semi-matroids also involve functional dependencies within $N$ (see the remark concluding Section 2.1). But otherwise semi-matroids are defined analogously: they are induced by polymatroids through vanishing of the respective submodular difference expressions.
@@
3.3. Graphoids, compositional semi-graphoids, and compositional graphoids
A semi-graphoid is called compositional if it satisfies the reversed implication, and a compositional graphoid is a semi-graphoid satisfying both implications. It is evident that ${\mathcal{M}}$ is a graphoid if and only if its dual model ${\mathcal{M}}^{\star}$ is a compositional semi-graphoid.
The above definitions, thus, delimit the frame $\mathfrak{gr}$ of graphoids, the frame $\mathfrak{gr}^{\star}$ of compositional semi-graphoids, and the frame $\mathfrak{cgr}$ of compositional graphoids. By definition, $\mathfrak{cgr}\subseteq\mathfrak{gr},\mathfrak{gr}^{\star}\subseteq\mathfrak{sg}$. Our notation reflects the fact that $\mathfrak{gr}^{\star}$ can be interpreted as the dual frame to $\mathfrak{gr}$ because ${\mathcal{M}}\in\mathfrak{gr}^{\star}(N)\,\Leftrightarrow\,{\mathcal{M}}^{\star}\in\mathfrak{gr}(N)$ for any variable set $N$ and ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$. It is known that $\mathfrak{gr}\setminus\mathfrak{pr}\neq\emptyset$ and $\mathfrak{pr}\setminus\mathfrak{gr}\neq\emptyset$.
3.4. Graphical CI frames
The family of undirected graphical models is then
This defines a CI frame $\mathfrak{ug}$, which is closed under copying and marginalization, but not under intersection. Well-known facts are that $\mathfrak{ug}\subseteq\mathfrak{pr}\subseteq\mathfrak{sg}$ [17]. The $\mathfrak{ug}$-frame has a finite “axiomatization” [32].
A popular class of CI models studied in probabilistic reasoning are directed acyclic graphical models. These are induced by directed graphs without directed cycles through the respective $d$-separation criterion [32]. The corresponding CI frame is closed under copying but not under marginalization [37, Example 3.1].
3.5. Other examples of abstract CI frames
Further options were pinpointed by one of the reviewers. One can consider the matroidal frame, a subframe of $\mathfrak{st}$ consisting of CI models induced by rank functions of matroids. The other option is the linear frame, consisting of CI models induced by linear polymatroids. Yet another option is the almost entropic frame, discussed in more details in Section 8.
@@ 4. Adhesion and self-adhesion
Matúš’s concept of self-adhesion can be viewed as an abstraction of the procedure used to derive non-Shannon information inequalities [43], which is commonly named the “copy lemma" method [12] in information-theoretical community. More details on the connection to this field are discussed in Section 8.
It has been observed recently in [5] that these concepts can also be defined within a simpler discrete environment of CI structures. We follow this link here. An auxiliary concept of consonant models is needed.
@@ 4.1. Self-adhesion
The self-adhesion of a CI model ${\mathcal{M}}$ is nothing but the adhesion of ${\mathcal{M}}$ with its own $\varphi$-copy, where $\varphi$ is an appropriate bijection. The derived concept of self-adhesivity is relative to a CI frame.
@@
  • Let $\mathfrak{f}$ be a CI frame closed under copying and marginalization. Given ${\mathcal{M}}\in\mathfrak{f}(N)$ and $L\subseteq N$ we say that ${\mathcal{M}}$ is self-adhesive at $L$ relative to $\mathfrak{f}$ if, for every bijection $\varphi:N\to M$ such that $L=N\cap M$ and $\varphi_{\,|L}$ is the identity, an adhesion of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ exists which is in the family $\mathfrak{f}(NM)$:
@@
    We say that ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive relative to $\mathfrak{f}$ if it is self-adhesive at every $L\subseteq N$.
  • The symbol $\mathfrak{f}^{\diamond}(N)$ will denote the family of self-adhesive models ${\mathcal{M}}\in\mathfrak{f}(N)$ relative to $\mathfrak{f}$. If $\mathfrak{f}(N)=\mathfrak{f}^{\diamond}(N)$ then we say that $\mathfrak{f}(N)$ is self-adhesive. The frame $\mathfrak{f}$ is self-adhesive if this holds for every finite $N$.
@@
The assumptions on $\mathfrak{f}$ (copying and marginalization) are needed to define the concept of self-adhesion. Clearly, a model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L\subseteq N$ relative to $\mathfrak{f}$ if and only if $\mathfrak{f}^{\diamond}({\mathcal{M}}|L)={\mathcal{M}}$.
@@
    This is to explain why the self-adhesivity concept is relative to a CI frame. Having two frames $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}$, imagine one finds a model ${\mathcal{M}}\in\mathfrak{f}_{1}(N)$ and a bijection $\varphi:N\to M$ with $L=N\cap M$ such that a self-adhesion ${\mathcal{E}}\in\mathfrak{f}_{2}(NM)$ of ${\mathcal{M}}$ for $\varphi$ exists while there is no such a self-adhesion within $\mathfrak{f}_{1}(NM)$. Then ${\mathcal{M}}$ is self-adhesive at $L$ relative to $\mathfrak{f}_{2}$ but not relative to $\mathfrak{f}_{1}$ (see Example 4.4 below).
@@
    It follows from the definition that, given ${\mathcal{M}}\in\mathfrak{f}(N)$, the question whether ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ can be answered on basis of the characterization/description of families $\mathfrak{f}(M)$, where $|N|\leq|M|\leq 2\cdot|N|$. Nevertheless, further reduction is possible. For example, it is immediate that any ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L:=N$. Additional reductions, based on further assumptions on the frame, are discussed below.
Example 4.4.
( self-adhesivity is relative to a frame )
Consider frames $\mathfrak{f}_{1}:=\mathfrak{ug}$, $\mathfrak{f}_{2}:=\mathfrak{sg}$ and the variable set $N:=\{a,b,c\}$. The model :=${\mathcal{M}}:=\{\,bc|a\,\}\subseteq{\mathcal{K}}(N)$ belongs to $\mathfrak{ug}(N)$: it is induced by the graph $G$ over $N$ with two edges, those connecting $a$ with the other two nodes. Then ${\mathcal{M}}\in\mathfrak{sg}^{\diamond}(N)$ but ${\mathcal{M}}\not\in\mathfrak{ug}^{\diamond}(N)$. To this end consider the set $L:=bc$ and a bijection $\varphi:N\to M:=\{\bar{a},b,c\}$ with $\varphi(a):=\bar{a}$, $\varphi(b):=b$, and $\varphi(c):=c$ (see Figure 2). Then ${\mathcal{M}}$ is not self-adhesive at $L$ relative to $\mathfrak{ug}$. Indeed, a potential adhesion ${\mathcal{E}}$ of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ in $\mathfrak{ug}(NM)$ has to satisfy
Figure 2. The bijection $\varphi:N\to M$ considered in Examples 4.4 and 4.6.
If a CI frame $\mathfrak{f}$ is closed under intersection then the self-adhesive closure of any model at any set is easy to compute by applying the $\mathfrak{f}$-closure. In particular, the question of self-adhesivity testing can be simplified.
@@
Let $\mathfrak{f}$ be a CI frame closed under copying, marginalization, intersection, and ascetic extension. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $L\subseteq N$, the self-adhesive closure ${\mathcal{R}}=\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$ exists and can be computed as follows. Having fixed a bijection $\varphi:N\to M$ with $L=N\cap M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$, apply the next steps:
@@
    apply the $\mathfrak{f}$-closure to it: ~$\widetilde{{\mathcal{E}}}:=\mathfrak{f}({\mathcal{E}})$,  (see Definition 2.20)